Consider a small economy where the total population is 20,000. The size of the labour force is 10,000, while the number of people employed is 7,000. The unemployment rate in this economy is 15%. (2 marks) N d) In 2017 the GDP of Ghana in current prices was GHC 205914 million, while the GDP in constant price was GHC 39175.04 million. The GDP-deflator for 2017 is therefore is 425.63 %. 


1
Expert's answer
2021-09-20T12:01:58-0400

unemployment rate = "\frac{Unemployed people}{Total labor force} \times 100"

Unemployed people = 10000-7000

= 3000


unemployment rate = "\frac{3000}{10000}\times100"

unemployment rate as per the data given = 30% Therefore the rate given in the question is basically wrong.


GDP Deflator = "\frac{Nominal GDP}{Real GDP}\times 100"


GDP Deflator = "\frac{205914 million}{39175.04 million}\times 100"


GDP Deflator = 525.63% Again the GDP Deflator indicated in the question is wrong.